Grade 11Mechanics

A room has dimension 3m×4m×5m.A fly starting at one corner end up at the dramatically opposite corner what is the magnitude of its displacement?if fly walk,what is the length of the shortest path

Jitendra pandey
8 Years agoGrade 11

1 Answer

Anu
7 Years ago
As a fly starting at one corner ends up at the diametrically opposite corner then we have to use the 3D version of Pythagoras which is d^2= a^2 + b^2+ c^2=3^2+4^2+5^2=9+16+25=50 then d^2=50 or d=5√2 . So 5√2 is the correct answer . I hope this will help you